Solve x2 dy/dx = y2 - xy. given that y = 1 when x = 1


1
Expert's answer
2022-05-16T16:43:57-0400

Solution

Dividing by x2 we’ll get equation  dy/dx = y2 / x2  – y/x

Substitution    z=y/x, y=z*x  =>  dy/dx = x(dz/dx) + z

From given equation  x(dz/dx)  = z2 – 2z  =>  "\frac{dz}{z^2-2z}=\frac{dx}{x}"   =>  "\int\frac{dz}{z^2-2z}=\int\frac{dx}{x}"   =>   "\frac{1}{2}\int\left(\frac{1}{z-2}-\frac{1}{z}\right)dz=\int\frac{dx}{x}"   =>  ln|z–2| – ln|z| = 2ln|x| + C

If y = 1 when x = 1 then z = 1 when x = 1. Substitution this into last equality gives 0 – 0 = 2*0 + C  =>  C = 0.

So solution of the given equation is

ln|z–2| – ln|z| = 2ln|x|  => ln(|z–2| / |z|) = 2ln|x|  => |(y-2x)/y| = x2 , 2x – y = y x2 => y(1 + x2) = 2x => y(x) = 2x/(1 + x2)     

Answer

y(x) = 2x/(1 + x2)